In a binomial distribution, n = 5 and π=0.32π=0.32 . Find the probabilities of the following events. (Round your answers to 4 decimal places.)

a. x=2x=2
Answer #1

Solution :

Given that ,

p = 0.32

1 - p = 1 -0.32 = 0.68  

n = 5

X = 2

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * (1 - p)n - x

P(X = 2) = ( (5! / (5 - 2)!) * 0.322 * 0.685 - 2

= 0.3220

Probability = 0.3220